OpenStudy (anonymous):

A fisherman decided to stock his favorite lake with a few fish that are not native to the region. The fish he introduces have no natural predators in the lake and they experience a large exponential growth. The introduction of an invasive species can be very detrimental to an ecosystem and there have already been noted effects. The National Fish and Wildlife Foundation has asked for your help. Create the data to fill in the population table. Be creative when selecting the P(x) values, but remember that this is an exponential growth model. @SolomonZelman

OpenStudy (anonymous):

Notice that you even need to come up with how many fish the fisherman introduces to the lake. Months (x) Population P(x) 0 1 2 3

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

okay so you want to create the table, say the fish population doubles every month

OpenStudy (anonymous):

ok so times 2

ganeshie8 (ganeshie8):

yup ! and wat about the initial starting population lets say 100 ?

OpenStudy (anonymous):

100 200 400 800

ganeshie8 (ganeshie8):

then, after 1 month, the population doubles, P(x) = 100*2 = 200

ganeshie8 (ganeshie8):

Yes ! good

OpenStudy (anonymous):

we only have to put for three months right?

ganeshie8 (ganeshie8):

0 100*2^0 = 100 1 100*2^1 = 200 2 100*2^2 = 400 3 100*2^3 = 800 .... x 100*2^x

ganeshie8 (ganeshie8):

correct, the table asks oly for the values upto 3 months

OpenStudy (anonymous):

ok thats it?

ganeshie8 (ganeshie8):

thats it :)

OpenStudy (anonymous):

ok lol there is another part to this. Describe how to create the exponential function represented by your table. Use complete sentences and arrive at a final exponential function, P(x).

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

we're done wid that part also :) you need to describe the function though

ganeshie8 (ganeshie8):

exponential function that represents the table is below :- \(\large P(x) = 100*2^x\)

OpenStudy (anonymous):

how do i describe it

ganeshie8 (ganeshie8):

describe wat each part describes above, in ur own words

ganeshie8 (ganeshie8):

100 = starting population of fish right ? 2 = rate of growth of fish per month ? x = months passed ?

ganeshie8 (ganeshie8):

frame that info into a short paragraph..

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

let's say i use this data. how would you calculate when the population of the fish reaches 500? @ganeshie8

OpenStudy (anonymous):

and how would you find the rate of growth per week?